OpenStudy (anonymous):

What would be the extension of a system of a spring and a mass, 'M' where in the spring itself has a mass, 'm'?

OpenStudy (primeralph):

It's complex to include the mass of the spring because it expands sectionally. In most cases, only include the mass of the object. So M*g = k*x g is accel due to gravity, x is the extension and k is the force constant of the spring. So, the extension x = M*g/k.

OpenStudy (rajat97):

i think we need to go to the centre of mass frame i'm not sure but i'll try and post it:)

OpenStudy (fifciol):

go ahead @rajat97 :)

OpenStudy (rajat97):

yup i'm trying but it's midnight (IST) so i'm feeling a bit sleepy but i'm trying it:) and what do you think will it be done by the c-frame??

OpenStudy (fifciol):

yes

OpenStudy (vincent-lyon.fr):

You need to write net force = 0 for an infinitesimal part of the spring of mass dm.

OpenStudy (rajat97):

i'll post a drawing of what i thought but i doubt it as the displacement of the centre of mass of the spring may not be equal to the displacement of the c.o.m. of the system

OpenStudy (rajat97):

i think we need to go with integration as @Vincent-Lyon.Fr said:)

OpenStudy (rajat97):

every infinitesimal part of the spring will experience two forces on is the restoring spring force and other is the weight of the block and that of itself

OpenStudy (rajat97):

is it right?

OpenStudy (rajat97):

but how will we relate it to the distance??

OpenStudy (rajat97):

if i go by the c.o.m. frame , i get the extension=(M+m)g/k where k is the spring constant

OpenStudy (rajat97):

what do you think??

OpenStudy (vincent-lyon.fr):

No, you are studying a static problem. There is no different point of view in different frames.

OpenStudy (rajat97):

oh! thanks for that:) but i don't know what a static problem really is

OpenStudy (vincent-lyon.fr):

Use elemental length (at rest) \(dl\) with mass \(dm=\mu \;dl\) and write that net force is zero, callinf \(\vec T(x)\) the tension of the upper part acting on the lower part. I have no time to derive it now, but I'll do it after dinner.

OpenStudy (vincent-lyon.fr):

If you have studied waves already, it is the same kind of demo as when you derive propagation equation on a string.

OpenStudy (rajat97):

yeah sure!

OpenStudy (rajat97):

i've not studied waves

OpenStudy (rajat97):

even i need to go now

OpenStudy (anonymous):

My recollection is that Prof. Lewin mentioned the answer to this, found by a colleague, during one of his mit.edu MIT lectures. One-third of the spring mass was added to the mass at the end of the spring, if I recall correctly.

OpenStudy (vincent-lyon.fr):

Let's consider a spring of length at rest L, mass m and constant k. According to the additivity of spring constants, any part of this spring of length dL will have a constant kL/dL. Let's write e = kL. e is the equivalent of Young's modulus for this spring. Let's call x the position of any coil of the spring when it is at rest and u(x), the displacement of this coil when the spring is under tension. Hooke's law becomes: T(x) = e du/dx or du/dx = T(x)/e = T(x)/kL (from now, we will not use e anymore) Then \(u(x)=\int_0^x du=\frac{1}{kL}\int_0^x T(x)dx\) Let's pull the spring with a given force F. When the spring has no mass, or when it has a mass but is horizontal, tension T is the same throughout the length of the spring and is equal to F.  Then u(x) is the area under the curve from 0 to x divided by kL. For x=L, total lengthening of the spring will be: \(\Delta L=u(L)=FL/kL=F/k\) which is the expected answer. Now, if the spring has a mass and is hung vertically, then tension will be greatest at the top and decreasing as you go from x=0 to L since it has to balance the weight of the bottom part of the spring. T(L) = F by choice of applying a force F there; T(0) = F + mg The graph looks like:  Now, the total area under the curve will be (F + mg/2)L and the new elongation will be : F/k + mg/2k You now see that the mass m of the springs acts as an extra m/2 mass acting on a massless spring. If F is the weight of a mass M, then extension will be (M + m/2)g/k , which answers cskc97's initial question.